OpenStudy (anonymous):

find all the solutions to secx=sqrt2

OpenStudy (gorv):

secx=sqrt2 secx=1/cosx 1/cosx=sqrt2 1/sqrt2=cosx at x= 45or pi/4 cos give this value

geerky42 (geerky42):

"find all the solutions" So \(\pi/4\) is not only solution here. In range \([-\pi,\pi)\), \(\cos x = \dfrac{1}{\sqrt2}\) is true at \(x=\pi/4\) and \(\)\(-\pi/4\) So all solutions are \(x = \dfrac{\pi}{4}+2\pi~n,~ n\in\mathbb{Z}\) and \( -\dfrac{\pi}{4}+2\pi~n,~ n\in\mathbb{Z}\)

OpenStudy (gorv):

that is understandable